# WaveGuard: Robust Deepfake Detection and Source Tracing via Dual-Tree Complex Wavelet and Graph Neural Networks

### v3

@@ Abstract
Deepfake technology has great potential in the field of media and entertainment, but it also brings serious risks, including privacy disclosure and identity fraud. To counter these threats, proactive forensic methods have become a research hotspot by embedding invisible watermark signals to build active protection schemes. However, existing methods are vulnerable to watermark destruction under malicious distortions, which leads to insufficient robustness. Moreover, embedding strong signals may degrade image quality, making it challenging to balance robustness and imperceptibility. To address these issues, this paper proposes a proactive watermarking framework called WaveGuard, which explores frequency domain embedding and graph-based structural consistency optimization. In this framework, the watermark is embedded into the high-frequency sub-bands by dual-tree complex wavelet transform (DT-CWT) to enhance the robustness against distortions and deepfake forgeries. By leveraging joint sub-band correlations and utilizing specific sub-band combinations, the framework performs robust extraction for source tracing and semi-robust extraction for deepfake detection. In addition, the framework adopts dense connectivity strategies to promote feature reuse and improve robustness under complex attacks. To enhance imperceptibility, we propose a Structural Consistency Graph Neural Network (SC-GNN), which constructs graph representations to extract structural features from the original and watermarked images, ensuring consistency between them and reducing perceptual artifacts to improve visual quality. Experimental results show that the proposed method performs exceptionally well in face swap and face replay tasks. The code has been published at https://github.com/vpsg-research/WaveGuard.
Index Terms:
Deepfake Detection, Source Tracing, Frequency-domain Embedding, Graph Neural Network (GNN).
@@ I Introduction
The rapid development of artificial intelligence generated content (AIGC) has greatly improved the ability of automation and innovation. However, its wide application also brings risks, especially the deepfake technology, which can generate forged media, threaten privacy, encourage identity theft and spread false information[1]. Deepfake generation techniques mainly includes face swap and face replay. Face swap[2] replaces the identity features of the target into the source image while preserving the background and semantic information. Face replay[3, 4] manipulates expressions or poses to map dynamic features onto the target image while maintaining identity consistency. Although deepfake has positive application in some fields, its abuse has caused serious social concern. Thus, it is very important to develop effective detection methods.
Currently, most deepfake detection methods rely on passive detection, which determines the authenticity of images by analyzing anomalous features present in forged content. Typical methods include spatial-domain artifact detection, such as color inconsistency, abnormal boundary mixing and irregular facial muscle movement[5, 6]. Other approaches explore frequency-domain features, identifying forgery clues like spectral distortions and encoding residuals[7]. In addition, some work also introduces biometric consistency (such as blink frequency and heartbeat signal) as auxiliary information to enhance the identification ability of forged content[8]. Although these method are effective in content authenticity verification, it usually comes into play after the generation and dissemination of forged content, and lacks effective traceability and interpretability.
@@
Recently, the concept of proactive approach has emerged as a preventive strategy, and its core idea is to embed invisible signals when the original content is released. This approach is mainly divided into proactive distortions[9, 10, 11, 12, 13] and proactive watermarks[14, 15, 16, 17, 18, 19, 20, 21, 22]. The former interferes with the generation process of a deepfake model by embedding invisible adversarial perturbation in the original image, destroying or counteracting the effect of the deepfake model, and making the generated content appear visually abnormal. While this provides early protection, it lacks traceability and verification capabilities. In addition, proactive distortions may also weaken the usability of deepfake technology in legitimate applications. In contrast, the proactive watermarks method is to embed the watermark information in the original content in advance, and then realize the dual functions of traceability and detectability by detecting the absence or change of the watermark. This method ensures copyright protection and source tracing, especially avoids the unexplained problems in the passive deepfake detection results. However, the existing proactive watermarks methods face challenges in robustness, particularly when dealing with complex deepfake attacks. As deepfake technology advances, the forgery techniques have become more subtle and sophisticated, making it increasingly difficult to extract watermarks effectively. Meanwhile, maintaining invisibility is equally important. Improving robustness often requires stronger embedding, which may introduce perceptual artifacts. Conversely, enhancing invisibility will reduce the robustness against distortion. Thus, balancing robustness and invisibility has become a key challenge for proactive watermarks.
@@
The remainder of this paper is organized as follows. Section II reviews the related work on digital watermarking and deepfake detection. Section III presents the proposed WaveGuard framework in detail. Section IV describes the experimental setup and reports the evaluation results. Section V concludes the paper.
@@ II-B Proactive Watermarks
Digital watermarking is a technology to protect copyrights, verify integrity, or source tracing by embedding hidden information in digital media. HiDDeN[26] is the first end-to-end trainable digital robust watermarking framework. It uses the encoder-decoder structure of deep neural networks to hide information and achieves efficient recovery under various image distortions, opening a new research direction in the field of digital watermarking. Aiming at the problem of non-differentiable JPEG compression, MBRS[27] further enhances the robustness of the model by randomly switching real JPEG, simulated JPEG, and noise-free layers in small batches. Similarly, CIN[28] integrates a reversible network and an attention module to enhance robustness while maintaining high invisibility and resistance to distortions. To overcome the limitations of traditional single-stage end-to-end frameworks, TSDL[29] employs a two-stage design to strengthen watermark robustness and improve model stability and convergence speed. Tang et al.[30]proposed a two-stage reversible watermarking framework that embeds watermarks into pseudo-Zernike moments using adaptive normalization and optimized DC-QIM, achieving strong resistance to distortions while enabling exact image recovery under benign conditions. Similarly, Huang et al.[31] proposed a texture-aware watermarking method that embeds in high-texture regions using an SSIM-guided regression model, achieving robustness to various attacks while maintaining perceptual quality. In addition, Wang et al.[32] proposed a watermarking framework using DCT-SVD embedding and a template-enhanced network to achieve robust synchronization and accurate extraction under geometric distortions. More recently, Zhu et al.[33] proposed a deep watermarking framework combining Lite-HRNet and a dual U-Net encoder, achieving robustness to screen-shooting and printing while preserving imperceptibility via adaptive noise and adversarial training. For cross-media distortion (such as print scanning, print camera and screen camera, etc.), researchers have proposed various solutions, including LFM[34], Stegstamp [35], RIHOOP [36], PIMoG [37], GDS-WM [38], and WRAP[39]. These methods target different distortion types by designing optimized strategies to ensure stable watermark extraction and verification after physical printing and re-digitization.
@@
In the emerging field of proactive face forensics, existing methods are limited. Yu et al. [14, 15] proposed an artificial fingerprint embedded within generative models for deepfake detection and source tracing, demonstrating strong scalability and robustness against complex noise. FaceSigns[16] differs from traditional robust watermarking by emphasizing robustness under common distortions while selectively failing under malicious distortions, thus achieving more accurate deepfake detection. Similarly, FaceGuard[17] achieves proactive deepfake detection by embedding semi-robust watermarks. Building on these efforts, Sepmark[18] proposed a separate watermarking framework that effectively addresses both common and malicious distortions by combining robust and semi-robust watermarking using a single encoder and separable decoder. At the same time, Wu et al.[19] proposed a new framework that adjusts robust watermarks into adversarial ones to improve detection and tracing capabilities. It also supports passive detection, making it easier to combine proactive and passive detection. Wang et al. [20] proposed an identity-perceptual watermarking framework that enables unified deepfake detection and source tracing by embedding content-dependent semantic watermarks. In addition, Lampmark[21] uses an untrained pre-defined pattern watermark to actively defend against deepfake, combining structural sensitivity and robustness to achieve efficient detection. With the growing use of multifunctional watermarking, Dual Defense[22] employs an original-domain feature impersonation strategy to jointly embed adversarial and traceable watermarks, thereby actively disrupting face swap generation and enabling reliable traceability.
Overall, although various methods have explored watermark robustness and deepfake detection, there are still challenges in improving robustness and imperceptibility. Some methods rely on spatial domain embedding and are easily affected by tampering, while others lack effective visual consistency maintenance mechanism. In contrast, our method integrates frequency domain embedding and structural consistency optimization, and achieves the unity of forgery detection and source tracking while taking into account robustness and invisibility.
@@ III-A1 Dual-tree complex wavelet transform
III-A2 Joint Sub-bands
Joint sub-band correlation was first proposed by Huan et al.[40], who significantly enhanced the robustness of video watermarking through the directional selectivity and translation invariance of DT-CWT. To further verify its role in image watermark embedding, we designed a simple experiment. A standard grayscale image (256 × 256 pixels) was decomposed by DT-CWT to obtain a low-frequency sub-band and six high-frequency sub-bands. Watermarks were embedded in sub-band 1 and sub-band 3. The results show that, although the watermark is only embedded in sub-band 1 and sub-band 3, sub-band 4 and sub-band 6 carry some watermark information due to the redundancy between sub-bands. The correlation with the original watermark is significantly high. Fig. 3 illustrates the distribution of the watermark in the frequency domain: pink represents the watermark embedding area, blue indicates the propagation area, and arrows show the propagation path. There is natural redundancy between the joint sub-bands, which enables the watermark information to spread in multiple high-frequency sub-bands, so it still has strong recovery ability under local tampering and frequency disturbance. Based on the correlation design of joint sub-bands, watermark information can spread among multiple sub-bands, thus maintaining stable extraction in different attack scenarios, which has become the key supporting mechanism for the system to achieve robust traceability and semi-robust deepfake detection.
@@ III-B Model Architecture
Fig. 4: Overall Architecture of WaveGuard. Frequency Subband Selection Module ($FSSM$) extracts and processes high-frequency sub-bands in the frequency domain, providing essential features for watermark embedding, tracing, and detection. Attention Module ($AM$) highlights salient regions in the image and guides the features to focus on key areas, thereby enhancing the robustness and effectiveness of watermark embedding. The encoder ($En$) embeds the watermark into the high-frequency sub-bands and reconstructs the watermarked image $I_{w}$. The Noise Pool Layer ($NPL$) applies various distortions to $I_{w}$ to generate distorted images $I_{n}$. The tracer decoder ($Tr$) extracts the watermark $w_{t}$ from $I_{n}$ to evaluate its robustness, while the decoder ($De$) extracts the correct watermark $w_{d}$ under common distortions and outputs random-like messages under malicious distortions. In addition, Structural Consistency Graph Neural Network ($SC-GNN$) extracts the structural features of the original image $I_{o}$ and the watermarked image $I_{w}$ by constructing graph representation to ensure the consistency between them.
@@ III-B3 Encoder (En)
In Formula (1), $U$ represents the original high-frequency sub-band, $\alpha$ controls the watermark embedding strength, M is the mask generated by the attention mechanism to weight the watermark embedding area, and $\text{wm}_{\text{diff}}$ is the generated watermark difference map. Through this process, the watermark information is embedded into the original high-frequency sub-band, forming a watermarked high-frequency sub-band $U_{\text{embed}}$. The post-processing module uses the inverse dual-tree complex wavelet transform to reconstruct the watermarked U-channel, merges it with the Y and V channels to form watermarked YUV images, and converts them to RGB images $I_{w}$, completing the embedding process.
@@ III-B5 Separable Decoder
The two decoders adopt a DenseNet structure similar to that of the encoder. The decoder takes the high-frequency sub-band embedded with the watermark as input, generates the initial feature map through the feature extraction module, and gradually fuses the decoded information through the multi-layer DenseBlock module. Finally, the high-dimensional features are compressed by the down-sampling module to generate a single-channel watermark difference map, which is adjusted to the target size through convolution and interpolation, and finally restored to one-dimensional watermark information.
Tracer and Detector share the same structure in design but differ in sub-band selection and target tasks. To enhance robustness, Tracer uses the 1st, 3rd, 4th, and 6th high-frequency sub-bands to extract features and handle various distortions, achieving watermark traceability with the goal of $W_{t}=W_{o}$. The Detector decoder only uses the 1st and 3rd sub-bands to extract features for the semi-robust requirement of forgery and tampering detection. Its target is $W_{d}=W_{o}$ for normal distortions and $W_{d}\neq W_{o}$ for malicious distortions. This difference in sub-band selection ensures that the two decoders meet the requirements for robust extraction and efficient detection, respectively.
@@ III-B6 Structural Consistency Graph Neural Network (SC-GNN))
First, the original image and the watermarked image will be sent to the Graph Construction Module. In this module, the pixels of each image are regarded as nodes in the graph, and the edges of the graph are constructed by 4- neighborhood connection (upper, lower, left and right adjacent pixels), thus generating graph representations of the original image and the watermarked image. In the graph construction module, each pixel of the image corresponds to a node, and the connection relationship (edge) between nodes is defined by the spatial relationship between adjacent pixels. Then, the graph data is sent to the graph convolution network (GCN). By capturing the spatial dependence between nodes, GCN extracts global and local features, optimizes the structural consistency between the watermark image and the original image, and reduces perceptual artifacts, thus improving the invisibility and robustness of the watermark. In the training process, the structural consistency loss function is used to further ensure the structural consistency between the watermark and the original image.

### v4

@@ Abstract
Deepfake technology has great potential in the field of media and entertainment, but it also brings serious risks, including privacy disclosure and identity fraud. To counter these threats, proactive forensic methods have become a research hotspot by embedding invisible watermark signals to build active protection schemes. However, existing methods are vulnerable to watermark destruction under malicious distortions, which leads to insufficient robustness. Moreover, embedding strong signals may degrade image quality, making it challenging to balance robustness and imperceptibility. Although watermarked images look natural, their underlying structures are often different from the original images, which is ignored by traditional watermarking methods. To address these issues, this paper proposes a proactive watermarking framework called WaveGuard, which explores frequency domain embedding and graph-based structural consistency optimization. In this framework, the watermark is embedded into the high-frequency sub-bands by dual-tree complex wavelet transform (DT-CWT) to enhance the robustness against distortions and deepfake forgeries. By leveraging joint sub-band correlations and selected sub-band combinations, the framework enables robust source tracing and semi-robust deepfake detection. To enhance imperceptibility, we propose a Structural Consistency Graph Neural Network (SC-GNN) that constructs graph representations of the original and watermarked images to ensure structural consistency and reduce perceptual artifacts. Experimental results show that the proposed method performs exceptionally well in face swap and face replay tasks. The code has been published at https://github.com/vpsg-research/WaveGuard.
@@ I Introduction
The rapid development of artificial intelligence generated content (AIGC) has greatly improved the ability of automation and innovation. However, its wide application also brings risks, especially the deepfake technology, which can generate forged media, threaten privacy, encourage identity theft and spread false information[1]. Deepfake generation techniques mainly includes face swap and face replay. Face swap[2], [3], [4], [5], [6], replaces the identity features of the target into the source image while preserving the background and semantic information. Face replay[7, 8] manipulates expressions or poses to map dynamic features onto the target image while maintaining identity consistency. Although deepfake has positive application in some fields, its abuse has caused serious social concern. Thus, it is very important to develop effective detection methods.
Currently, most deepfake detection methods rely on passive detection, which determines the authenticity of images by analyzing anomalous features present in forged content. Typically, quantization-based three-dimensional convolutional neural networks (3D CNNs) are employed to enhance the discrimination of spatiotemporal features [9]. Other studies use perturbation-domain alignment to improve cross-domain robustness [10]. In addition, researchers have explored frequency-domain analysis to detect spectral distortions and encoding residuals [11]. Lightweight video feature modeling methods have also been proposed to boost detection efficiency while maintaining high accuracy [12]. Although these methods are effective in content authenticity verification, it usually comes into play after the generation and dissemination of forged content, and lacks effective traceability and interpretability.
@@
Recently, the concept of proactive approach has emerged as a preventive strategy, and its core idea is to embed invisible signals when the original content is released. This approach is mainly divided into proactive distortions[13], [14], [15], [16], [17] and proactive watermarks[18], [19], [20], [21], [22], [23], [24], [25], [26]. The former interferes with the generation process of a deepfake model by embedding invisible adversarial perturbation in the original image, destroying or counteracting the effect of the deepfake model, and making the generated content appear visually abnormal. While this provides early protection, it lacks traceability and verification capabilities. In addition, proactive distortions may also weaken the usability of deepfake technology in legitimate applications. In contrast, the proactive watermarks method is to embed the watermark information in the original content in advance, and then realize the dual functions of traceability and detectability by detecting the absence or change of the watermark. This method ensures copyright protection and source tracing, especially avoids the unexplained problems in the passive deepfake detection results. However, the existing proactive watermarks methods face challenges in robustness, particularly when dealing with complex deepfake attacks. As deepfake technology advances, the forgery techniques have become more subtle and sophisticated, making it increasingly difficult to extract watermarks effectively. Meanwhile, maintaining invisibility is equally important. Improving robustness often requires stronger embedding, which may introduce perceptual artifacts. Conversely, enhancing invisibility will reduce the robustness against distortion. Thus, balancing robustness and invisibility has become a key challenge for proactive watermarks.
@@
The remainder of this paper is organized as follows. Section II reviews the related work on proactive defense, digital watermarking, and proactive forensics. Section III presents the proposed WaveGuard framework in detail. Section IV describes the experimental setup and reports the evaluation results. Section V concludes the paper.
@@ II-B Proactive Watermarks
Digital watermarking is a technology to protect copyrights, verify integrity, or source tracing by embedding hidden information in digital media. HiDDeN[30] is the first end-to-end trainable digital robust watermarking framework. It uses the encoder-decoder structure of deep neural networks to hide information and achieves efficient recovery under various image distortions, opening a new research direction in the field of digital watermarking. To address challenges introduced by JPEG compression, MBRS[31] introduced a robustness-oriented approach by dynamically switching among different encoding settings, including real and simulated JPEG as well as noise-free pathways, thereby enhancing adaptability. Similarly, CIN[32] further advanced robustness and invisibility by combining a reversible transform with attention mechanisms. In addition, TSDL[33] replaced the conventional single-stage architecture with a sequential dual-phase model, resulting in more stable embedding and faster convergence during training. Tang et al.[34]proposed a two-stage reversible watermarking framework that embeds watermarks into pseudo-Zernike moments using adaptive normalization and optimized DC-QIM, achieving strong resistance to distortions while enabling exact image recovery under benign conditions. ARWGAN[35] introduces an attention-guided GAN-based watermarking model that enhances robustness through adversarial training under diverse distortions. Similarly, Huang et al.[36] proposed a texture-aware watermarking method that embeds in high-texture regions using an SSIM-guided regression model, achieving robustness to various attacks while maintaining perceptual quality. In addition, Wang et al.[37] proposed a watermarking framework using DCT-SVD embedding and a template-enhanced network to achieve robust synchronization and accurate extraction under geometric distortions. More recently, Zhu et al.[38] proposed a deep watermarking framework combining Lite-HRNet and a dual U-Net encoder, achieving robustness to screen-shooting and printing while preserving imperceptibility via adaptive noise and adversarial training. For cross-media distortion (such as print scanning, print camera and screen camera, etc.), researchers have proposed various solutions, including LFM[39], Stegstamp [40], RIHOOP [41], PIMoG [42], GDS-WM [43], and WRAP[44]. These methods target different distortion types by designing optimized strategies to ensure stable watermark extraction and verification after physical printing and re-digitization.
@@
In the emerging field of proactive face forensics, existing methods are limited. Yu et al. [18, 19] proposed an artificial fingerprint embedded within generative models for deepfake detection and source tracing, demonstrating strong scalability and robustness against complex noise. FaceSigns[20] differs from traditional robust watermarking by emphasizing robustness under common distortions while selectively failing under malicious distortions, thus achieving more accurate deepfake detection. Similarly, FaceGuard[21] achieves proactive deepfake detection by embedding semi-robust watermarks. Building on these efforts, Sepmark[22] proposed a separate watermarking framework that effectively addresses both common and malicious distortions by combining robust and semi-robust watermarking using a single encoder and separable decoder. At the same time, Wu et al.[23] proposed a new framework that adjusts robust watermarks into adversarial ones to improve detection and tracing capabilities. It also supports passive detection, making it easier to combine proactive and passive detection. Wang et al. [24] proposed an identity-perceptual watermarking framework that enables unified deepfake detection and source tracing by embedding content-dependent semantic watermarks. EditGuard[45] proposes a proactive watermarking framework that enables zero-shot tamper localization and copyright protection via dual image-bit steganography. In addition, Lampmark[25] uses an untrained pre-defined pattern watermark to actively defend against deepfake, combining structural sensitivity and robustness to achieve efficient detection. With the growing use of multifunctional watermarking, Dual Defense[26] employs an original-domain feature impersonation strategy to jointly embed adversarial and traceable watermarks, thereby actively disrupting face swap generation and enabling reliable traceability.
Overall, although the existing watermarking methods show some robustness under normal distortion conditions, they often fail in the face of advanced deepfake operations. This kind of attack will reconstruct the semantic content and structural layout of the image, and the deepfake model usually re-renders the face region in a generative way. However, most watermarking methods only embed low-level signals in pixel space or traditional frequency domain, which lacks the adaptability to semantic content changes. Meanwhile, the deepfake model has the ability of end-to-end optimization, which can automatically suppress or erase the imperceptible disturbance signal during the generation process, resulting in the watermark information being overwritten. In this paper, WaveGuard embeds watermarks into directional frequency sub-bands and incorporates graph neural networks to reinforce structural consistency. Although it does not explicitly model semantic content, this structure-aware design exhibits strong resilience to distortions induced by semantic-level manipulations (e.g., generative face reconstruction), enabling reliable extraction while maintaining a favorable trade-off between robustness and invisibility.
@@ III-A1 Dual-tree complex wavelet transform
III-A2 Joint Sub-bands
Joint sub-band correlation was first proposed by Huan et al.[46], who significantly enhanced the robustness of video watermarking through the directional selectivity and translation invariance of DT-CWT. To further verify its role in image watermark embedding, we designed a simple experiment.
@@ III-A2 Joint Sub-bands
A standard grayscale image (256 × 256 pixels) was decomposed by DT-CWT to obtain a low-frequency sub-band and six high-frequency sub-bands. Watermarks were embedded in sub-band 1 and sub-band 3. The results show that, although the watermark is only embedded in sub-band 1 and sub-band 3, sub-band 4 and sub-band 6 carry some watermark information due to the redundancy between sub-bands. The correlation with the original watermark is significantly high. Fig. 3 illustrates the distribution of the watermark in the frequency domain: pink represents the watermark embedding area, blue indicates the propagation area, and arrows show the propagation path. There is natural redundancy between the joint sub-bands, which enables the watermark information to spread in multiple high-frequency sub-bands, so it still has strong recovery ability under local tampering and frequency disturbance. Based on the correlation design of joint sub-bands, watermark information can spread among multiple sub-bands, thus maintaining stable extraction in different attack scenarios, which has become the key supporting mechanism for the system to achieve robust traceability and semi-robust deepfake detection.
@@ III-B Model Architecture
Fig. 4: Overall architecture of WaveGuard. Frequency Subband Selection Module (FSSMFSSM) extracts and processes high-frequency sub-bands in the frequency domain, providing essential features for watermark embedding, tracing, and detection. Attention Module (AMAM) highlights salient regions in the image and guides the features to focus on key areas, thereby enhancing the robustness and effectiveness of watermark embedding. The encoder (EnEn) embeds the watermark into the high-frequency sub-bands and reconstructs the watermarked image $I_{w}$. The Noise Pool Layer (NPLNPL) applies various distortions to $I_{w}$ to generate distorted images $I_{n}$. The tracer decoder (TrTr) extracts the watermark $w_{t}$ from $I_{n}$ to evaluate its robustness, while the decoder (DeDe) extracts the correct watermark $w_{d}$ under common distortions and outputs random-like messages under malicious distortions. In addition, Structural Consistency Graph Neural Network (SCGNNSC-GNN) extracts the structural features of the original image $I_{o}$ and the watermarked image $I_{w}$ by constructing graph representation to ensure the consistency between them.
@@ III-B3 Encoder (En)
In Formula (1), $U$ represents the original high-frequency sub-
band, $\alpha$ controls the watermark embedding strength, M is the mask generated by the attention mechanism to weight the watermark embedding area, and $\text{wm}_{\text{diff}}$ is the generated watermark difference map. Through this process, the watermark information is embedded into the original high-frequency sub-band, forming a watermarked high-frequency sub-band $U_{\text{embed}}$. The post-processing module uses the inverse dual-tree complex wavelet transform to reconstruct the watermarked U-channel, merges it with the Y and V channels to form watermarked YUV images, and converts them to RGB images $I_{w}$, completing the embedding process.
@@ III-B5 Separable Decoder
The two decoders adopt a DenseNet structure similar to that of the encoder. The decoder takes the high-frequency sub-band embedded with the watermark as input, generates the initial feature map through the feature extraction module, and gradually fuses the decoded information through the multi-layer DenseBlock module. Finally, the high-dimensional features are compressed by the down-sampling module to generate a single-channel watermark difference map, which is adjusted to the target size through convolution and interpolation, and finally restored to one-dimensional watermark information. Tracer and Detector share the same structure in design but differ in sub-band selection and target tasks. To enhance robustness, Tracer uses the 1st, 3rd, 4th, and 6th high-frequency sub-bands to extract features and handle various distortions, achieving watermark traceability with the goal of $W_{t}=W_{o}$. The Detector decoder only uses the 1st and 3rd sub-bands to extract features for the semi-robust requirement of forgery and tampering detection. Its target is $W_{d}=W_{o}$ for normal distortions and $W_{d}\neq W_{o}$ for malicious distortions. This difference in sub-band selection ensures that the two decoders meet the requirements for robust extraction and efficient detection, respectively.
@@ III-B6 Structural Consistency Graph Neural Network (SC-GNN))
Although the watermarked image may appear visually natural, its underlying structural distribution often deviates from the original image. Traditional watermarking methods typically overlook structural consistency, leading to local artifacts or edge distortions that degrade perceptual quality. To address this, we propose a Structural Consistency Graph Neural Network (SC-GNN).
Specifically, the $U$ channels of both the original and watermarked images are extracted as single-channel inputs and converted into sparse graphs based on 4-neighborhood connectivity (top, bottom, left, and right), where each pixel is treated as a node with its normalized intensity as the node feature. The constructed graph is fed into a two-layer Graph Convolutional Network (GCN): the first layer maps the node features to 16 dimensions with ReLU activation, while the second projects them to scalars. A global average pooling operation is then applied to obtain a graph-level embedding vector. This embedding vector is used to quantify the structural difference between the original and watermarked images and serves as the input to the structural consistency loss to preserve structural consistency.